OpenStudy (anonymous):

Find the dimensions of a rectangle whose perimeter is 22 meters and whose area is 28 square meters. The side of the retangle measure ? Meters

OpenStudy (campbell_st):

so you have for the perimeter where length = l and width = w 22 = 2l + 2w or 11 = l + w and for area 28 = lw take the perimeter equation and make w the subject and substitute into the area equation... then solve for l... when you get l, substitute into either equation to find w

OpenStudy (yttrium):

We know that LW = 28 and 2(L+W) = 22 You can actually solve this using substitution, considering LW = 28 L = 28/W Therefore, considering 2(L+W) = 22 \[2(\frac{ 28 }{ w}+w) = 22\] Therefore, upon simplifying you now solve for W. And to solve for L (just use the equation L = 28/W And that's it! :D

OpenStudy (anonymous):

Thank you both